# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 6. ${\mathfrak{F}}$ as equivariantization of ${\mathfrak{F}}^{b}$
In this section we will use Theorem 5.13 to show that the equivalence ${\mathfrak{F}}^{b}$ intertwines rational $B$-actions on both categories, so that ${\mathfrak{F}}$ is the equivalence of their equivariantizations. As an application, we show that ${\mathfrak{F}}$ intertwines reflection functors $\Theta^{r}_{s}$ and $\Theta^{c}_{s}$, for all $s\in{\mathrm{I}}_{\mathrm{af}}$. For simplicity, we only consider the non-deformed versions of ${\mathfrak{F}}^{b}$ and ${\mathfrak{F}}$.
@@ 6.1. Rational group action
$-\otimes-:\ {\mathcal{C}}\times\mathrm{Rep}(K)\rightarrow{\mathcal{C}}$
@@ 6.2. Coinduction functor
$\widehat{{\mathfrak{F}}}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}{\ \xrightarrow{\sim}\ }\widehat{{\mathcal{O}}}^{0}_{\mathbb{C}}.$
@@
${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}$
@@
$\widehat{{\mathfrak{F}}}^{n}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}.$
@@
which maps $\varphi$ on the LHS to the homomorphism $m_{1}\in M_{1}\mapsto\varphi(m_{1})(1)$, and $\phi$ on the RHS to the homomorphism $m_{1}\in M_{1}\mapsto\big{(}u\in{\mathbf{U}}^{\mathrm{hb}}_{\zeta}\mapsto\phi(um_{1})\big{)}$.
@@
with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
@@
$-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}:\ (\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}\ \rightleftarrows\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\ :\mathrm{coind}_{r}.$
@@
with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$. This follows from [4, Thm 2.8] applied to the sequence of Hopf algebras ${\mathbf{u}}^{\geq}_{\zeta}\hookrightarrow{\mathbf{U}}^{\geq}_{\zeta}\twoheadrightarrow U{\mathfrak{b}}$. ∎
@@
$\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B},\quad M\mapsto M\otimes{\mathbb{C}}[B].$
It is right adjoint to the forgetful functor ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$, and induces an equivalence
(6.3) $\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B})_{B}.$
@@ 6.3. $\widehat{{\mathfrak{F}}}$ as equivariantization of $\widehat{{\mathfrak{F}}}^{n}$
[…]
Note that ${\mathbb{V}}_{c}$, ${\mathbb{V}}_{r}$ and ${\mathfrak{G}}$ are compatible with $\mathrm{rep}(B)$-action. Note also that the $\mathrm{rep}(B)$-action preserves the subcategory of objects admitting Verma flags. Hence by Proposition 5.11, ${\mathfrak{F}}$ is compatible with $\mathrm{rep}(B)$-action on such subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. Since any object can be resolved by objects admitting Verma flags, such compatibility extends to the whole category. ∎
@@
The equivalence $\widehat{{\mathfrak{F}}}^{n}:{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$ is compatible with the natural rational $B$-action on both sides. Therefore, $\widehat{{\mathfrak{F}}}$ can be obtained as the $B$-equivariantization of $\widehat{{\mathfrak{F}}}^{n}$.
@@
[…]
@@
[…]
@@
${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0}$
@@ 7. Graded multiplicities
Recall that the tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, which makes ${\mathrm{A}}$ a ${\mathbb{C}}^{\times}$-equivariant algebra. Using the equivalence (5.6), we can define a graded version of ${\mathcal{O}}^{0}_{\mathbb{C}}$ via
${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\ :=\ \mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{\tilde{{\mathcal{N}}}}^{*}{\mathrm{A}})\ \simeq\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}},$
with a degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}^{0}_{\mathbb{C}}$. In this section, we compute the graded multiplicities of simple module in Verma module (resp. of Verma module in projective module) in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$.
@@ 7.1. Periodic and generic polynomials
Let $l(-):W_{\mathrm{ex}}\rightarrow{\mathbb{Z}}$ be the length function. The extended affine Hecke algebra ${\mathbb{H}}_{\mathrm{ex}}$ is the ${\mathbb{Z}}[v^{\pm 1}]$-algebra generated by $\{H_{x}\}_{x\in W_{\mathrm{ex}}}$ modulo the relations
$\begin{cases}(H_{s}+v)(H_{s}-v^{-1})=0&\text{for }s\in{\mathrm{I}}_{\mathrm{af}},\\ H_{x}H_{y}=H_{xy}&\text{if }l(x)+l(y)=l(xy).\end{cases}$
@@
Let ${\mathbb{P}}$ be the free ${\mathbb{Z}}[v^{\pm 1}]$-module with basis $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$. It is equipped with a right ${\mathbb{H}}_{\mathrm{ex}}$-module structure (called the periodic Hecke module) via
$H^{\frac{\infty}{2}}_{x}\cdot H_{s}=\begin{cases}H^{\frac{\infty}{2}}_{xs}&\text{if }x\leq^{\frac{\infty}{2}}xs,\\ H^{\frac{\infty}{2}}_{xs}+(v^{-1}-v)H^{\frac{\infty}{2}}_{x}&\text{if }xs\leq^{\frac{\infty}{2}}x,\end{cases}$
for any $x\in W_{\mathrm{ex}}$ and $s\in{\mathrm{I}}_{\mathrm{af}}$. For $\lambda\in\Lambda$ we set $e(\lambda):=\sum_{w\in W}v^{l(w)}\cdot H^{\frac{\infty}{2}}_{t(\lambda)w}$. Let ${\mathbb{P}}^{\circ}$ be the ${\mathbb{H}}_{\mathrm{ex}}$-submodule of ${\mathbb{P}}$ generated by $e(\lambda)$ with $\lambda\in\Lambda$.
@@
    On ${\mathbb{P}}^{\circ}$ there is a unique involution ${\mathbb{P}}^{\circ}\rightarrow{\mathbb{P}}^{\circ}$, $P\mapsto\overline{P}$ that is ${\mathbb{H}}_{\mathrm{ex}}$-skew-linear (i.e. compatible with the bar-involution) and satisfies $\overline{e(\lambda)}=e(\lambda)$ for each $\lambda\in\Lambda$.
@@
We express $\underline{H}^{\frac{\infty}{2}}_{x}=\sum_{y}p_{y,x}\cdot H^{\frac{\infty}{2}}_{y}$, then $p_{y,x}\in{\mathbb{Z}}[v]$ is called the periodic Kazhdan–Lusztig polynomial.
Let $\hat{{\mathbb{P}}}$ be the completion of ${\mathbb{P}}$ consisting of the formal series in $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). For $\eta\in\Lambda$, consider the ${\mathbb{Z}}[v^{\pm 1}]$-linear operator $\langle\eta\rangle$ on ${\mathbb{P}}$ by $\langle\eta\rangle(H^{\frac{\infty}{2}}_{x})=H^{\frac{\infty}{2}}_{t(\eta)x}$, for any $x\in W_{\mathrm{ex}}$. We define two families of polynomials $\{q_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ and $\{q^{\prime}_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ in ${\mathbb{Z}}[v]$ via the following equations in $\hat{{\mathbb{P}}}$ (for any $x\in W_{\mathrm{ex}}$)
(7.1) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+v^{2}\langle-\alpha\rangle+v^{4}\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q_{y,x}\cdot H^{\frac{\infty}{2}}_{y},$
(7.2) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q^{\prime}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}.$
Here $q_{y,x}$ is called the generic Kazhdan–Lusztig polynomial.
@@
The generic polynomials $\{q_{x,y}\}_{x,y}$ were originally defined as the unique family of polynomials such that the element $\sum_{y}(-1)^{l(x)+l(y)}\overline{q}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}\in\hat{{\mathbb{P}}}$ is contained in $H^{\frac{\infty}{2}}_{x}+\sum_{y}v^{-1}{\mathbb{Z}}[v^{-1}]\cdot H^{\frac{\infty}{2}}_{y}$ and is invariant under the involution, see [25]. Under this definition, the equality (7.1) is proved in [25, Thm 3.5], see also [39, Thm 6.3].

### v4

@@ 6. $\mathfrak{F}$} as equivariantization of ${\mathfrak{F}}^{b}$
In this section we will use Theorem 5.23 to show that the equivalence ${\mathfrak{F}}^{b}$ intertwines rational $B$-actions on both categories, so that $\mathfrak{F}$} is the equivalence of their equivariantizations. This essentially follows the idea in the work of Arkhipov–Gaitsgory [3]. As an application, we show that $\mathfrak{F}$} intertwines reflection functors $\Theta^{r}_{s}$ and $\Theta^{c}_{s}$, for all $s\in{\mathrm{I}}_{\mathrm{af}}$. For simplicity, we only consider the non-deformed versions of ${\mathfrak{F}}^{b}$ and $\mathfrak{F}$}.
@@ 6.1. Rational group action
$\otimes-:\ {\mathcal{C}}\times\mathrm{Rep}(K)\rightarrow{\mathcal{C}}$
@@ 6.2. Coinduction functor
$\widehat{{\mathfrak{F}}}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}{\ \xrightarrow{\sim}\ }\widehat{{\mathcal{O}}}^{0}_{\mathbb{C}}.$
@@
${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}$
@@
$\widehat{{\mathfrak{F}}}^{n}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}.$
@@
which maps $\varphi$ on the LHS to the homomorphism $m_{1}\in M_{1}\mapsto\varphi(m_{1})(1)$, and maps $\phi$ on the RHS to the homomorphism $m_{1}\in M_{1}\mapsto\big(u\in{\mathbf{U}}^{\mathrm{hb}}_{\zeta}\mapsto\phi(um_{1})\big)$.
@@
with inverse functor $\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
@@
$\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}:\ (\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}\ \rightleftarrows\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\ :\mathrm{coind}_{r}.$
@@
with inverse functor $\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$. This follows from [3, Thm 2.8] applied to the sequence of Hopf algebras ${\mathbf{u}}^{\geq}_{\zeta}\hookrightarrow{\mathbf{U}}^{\geq}_{\zeta}\twoheadrightarrow U{\mathfrak{b}}$. ∎
@@
$\mathrm{coind}_{c}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B},\quad M\mapsto M\otimes{\mathbb{C}}[B].$
It is right adjoint to the forgetful functor ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$, and induces an equivalence
(6.3) $\mathrm{coind}_{c}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B})_{B}.$
@@ 6.3. $\widehat{{\mathfrak{F}}}$ as equivariantization of $\widehat{{\mathfrak{F}}}^{n}$
${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$${\mathcal{O}}^{0}_{S}$$\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$${\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$$\scriptstyle{{\mathfrak{F}}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}}$$\scriptstyle{{\mathbb{V}}_{r}}$$\scriptstyle{{\mathfrak{G}}}\scriptstyle{\simeq}$
Note that ${\mathbb{V}}_{c}$, ${\mathbb{V}}_{r}$ and $\mathfrak{G}$} are compatible with $\mathrm{rep}(B)$-action. Note also that the $\mathrm{rep}(B)$-action preserves the subcategory of objects admitting Verma flags. Hence by Proposition 5.21, $\mathfrak{F}$} is compatible with $\mathrm{rep}(B)$-action on such subcategory in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. Since any object can be resolved by objects admitting Verma flags, such compatibility extends to the whole category. ∎
@@
The equivalence $\widehat{{\mathfrak{F}}}^{n}:{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$ is compatible with the natural rational $B$-action on both sides. Therefore, $\widehat{{\mathfrak{F}}}$ can be obtained as the $B$-equivariantization of $\widehat{{\mathfrak{F}}}^{n}$.
@@
${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}$$\widehat{{\mathcal{O}}}_{\mathbb{C}}^{0}.$$\scriptstyle{\widehat{{\mathfrak{F}}}^{n}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{coind}_{c}}$$\scriptstyle{\mathrm{coind}_{r}}$$\scriptstyle{\widehat{{\mathfrak{F}}}}\scriptstyle{\simeq}$
@@
(6.4) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}$$\widehat{{\mathcal{O}}}_{\mathbb{C}}^{0}$$({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod})^{B}$$({\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}})^{B},$$\scriptstyle{\widehat{{\mathfrak{F}}}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{(\widehat{{\mathfrak{F}}}^{n})^{B}}\scriptstyle{\simeq}$
@@
${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0}$
@@ 7. Graded multiplicities
[…]
${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\ :=\ \mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{\widetilde{{\mathcal{N}}}}^{*}{\mathsf{A}})\ \simeq\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}},$
with a degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}^{0}_{\mathbb{C}}$. In this section, we compute the graded multiplicity of simple modules in Verma modules in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, which turns out to be given by the generic Kazhdan–Lusztig polynomials defined in [24] (see also [28, 37]). We also obtain formulas for graded multiplicity of Verma module in projective module.
@@ 7.1. Periodic and generic polynomials
Let $\ell(-):W_{\mathrm{ex}}\rightarrow{\mathbb{Z}}$ be the length function. The extended affine Hecke algebra ${\mathbb{H}}_{\mathrm{ex}}$ is the ${\mathbb{Z}}[v^{\pm 1}]$-algebra generated by $\{H_{x}\}_{x\in W_{\mathrm{ex}}}$ modulo the relations
$\begin{cases}(H_{s}+v)(H_{s}-v^{-1})=0&\text{for }s\in{\mathrm{I}}_{\mathrm{af}},\\ H_{x}H_{y}=H_{xy}&\text{if }\ell(x)+\ell(y)=\ell(xy).\end{cases}$
@@
Recall the semi-infinite order $\leq^{\frac{\infty}{2}}$ in Definition 5.6. Let $\mathbb{P}$} be the free ${\mathbb{Z}}[v^{\pm 1}]$-module with basis $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$. It is equipped with a right ${\mathbb{H}}_{\mathrm{ex}}$-module structure (called the periodic Hecke module) via the following formulas for any $x\in W_{\mathrm{ex}}$, $s\in{\mathrm{I}}_{\mathrm{af}}$ and $\gamma\in\Lambda/{\mathrm{Q}}$,
$H^{\frac{\infty}{2}}_{x}\cdot H_{s}$ $=\begin{cases}H^{\frac{\infty}{2}}_{xs}&\text{if }x\leq^{\frac{\infty}{2}}xs,\\ H^{\frac{\infty}{2}}_{xs}+(v^{-1}-v)H^{\frac{\infty}{2}}_{x}&\text{if }xs\leq^{\frac{\infty}{2}}x;\end{cases}$
$H^{\frac{\infty}{2}}_{x}\cdot H_{\gamma}$ $=H^{\frac{\infty}{2}}_{x\gamma}.$
For $\lambda\in\Lambda$ we set $e(\lambda):=\sum_{w\in W}v^{\ell(w)}\cdot H^{\frac{\infty}{2}}_{t(\lambda)w}$. Let $\mathbb{P}$$^{\circ}$ be the ${\mathbb{H}}_{\mathrm{ex}}$-submodule of $\mathbb{P}$} generated by $e(\lambda)$ with $\lambda\in\Lambda$.
@@
    On $\mathbb{P}$$^{\circ}$ there is a unique involution $\mathbb{P}$$^{\circ}\rightarrow{\mathbb{P}}^{\circ}$, $P\mapsto\overline{P}$ that is ${\mathbb{H}}_{\mathrm{ex}}$-skew-linear (i.e. compatible with the bar-involution of ${\mathbb{H}}_{\mathrm{ex}}$) and satisfies $\overline{e(\lambda)}=e(\lambda)$ for each $\lambda\in\Lambda$.
@@
We express $\underline{H}^{\frac{\infty}{2}}_{x}=\sum_{y}p_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y}$, then $p_{y,x}(v)\in{\mathbb{Z}}[v]$ is called the periodic Kazhdan–Lusztig polynomial.
Let $\hat{{\mathbb{P}}}$ be the completion of $\mathbb{P}$} consisting of the formal series in $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). For $\eta\in\Lambda$, consider the ${\mathbb{Z}}[v^{\pm 1}]$-linear operator $\langle\eta\rangle$ on $\mathbb{P}$} by $\langle\eta\rangle(H^{\frac{\infty}{2}}_{x})=H^{\frac{\infty}{2}}_{t(\eta)x}$, for any $x\in W_{\mathrm{ex}}$. We define two families of polynomials $\{q_{y,x}(v)\}_{x,y\in W_{\mathrm{ex}}}$ and $\{q^{\prime}_{y,x}(v)\}_{x,y\in W_{\mathrm{ex}}}$ in ${\mathbb{Z}}[v]$ via the following equations in $\hat{{\mathbb{P}}}$ (for any $x\in W_{\mathrm{ex}}$)
(7.1) $\big(\prod_{\alpha\in\Phi^{+}}(1+v^{2}\langle-\alpha\rangle+v^{4}\langle-2\alpha\rangle+\cdots)\big)(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y},$
(7.2) $\big(\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big)(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q^{\prime}_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y}.$
Here $q_{y,x}(v)$ is called the generic Kazhdan–Lusztig polynomial. Note that $q_{y,x}(1)=q_{y,x}^{\prime}(1)$.
@@
  1. (1)
    In [37] the periodic (resp. generic) polynomials are labeled by alcoves: $p_{y,x}(v)$ (resp. $q_{y,x}(v)$) is denoted by $p_{yA_{\mathrm{fun}},xA_{\mathrm{fun}}}(v)$ (resp. by $q_{yA_{\mathrm{fun}},xA_{\mathrm{fun}}}(v)$) in the loc. cit..
  2. (2)
    The generic polynomials $\{q_{x,y}(v)\}_{x,y}$ are originally defined as the unique family of polynomials such that the element $\sum_{y}(-1)^{\ell(x)+\ell(y)}{q}_{y,x}(v^{-1})\cdot H^{\frac{\infty}{2}}_{y}\in\hat{{\mathbb{P}}}$ is contained in $H^{\frac{\infty}{2}}_{x}+\sum_{y}v^{-1}{\mathbb{Z}}[v^{-1}]\cdot H^{\frac{\infty}{2}}_{y}$ and is invariant under the involution on $\hat{{\mathbb{P}}}$ (obtained by extending the involution on $\mathbb{P}$$^{\circ}$), see [24]. Under this definition, the equality (7.1) is proved in [24, Thm 3.5], see also [37, Thm 6.3].